OpenStudy (anonymous):

Which property is shown below? -12r + (9s + 3t) = (-12r + 9s) + 3t Commutative Associative Identity Distributive

OpenStudy (midhun.madhu1987):

a + b = b + a ---- Commutative a + (b + c) = (a + b) + c - Associative a * (b + c) = a * b + a * c --- Distributive a + 0 = 0 + a - ---- Identity

OpenStudy (anonymous):

Commutative- a+b=b+a Associative- (a+b)+c=a+(b+c) Identity-a+0=a Distributive- a*(b+c)= ab+ac

OpenStudy (midhun.madhu1987):

So which option seems to be correct for u?

OpenStudy (anonymous):

You beat me -_- @midhun.madhu1987

OpenStudy (midhun.madhu1987):

lol :P

OpenStudy (anonymous):

Associative

OpenStudy (midhun.madhu1987):

Cool... u got it :)